OpenStudy (anonymous):

Find the exponential function y=Ce^kt that passes through the two points: (2,3/2) and (4,5)

OpenStudy (amistre64):

we know t and y, should we solve for k?

OpenStudy (anonymous):

K and C

OpenStudy (amistre64):

\[5=Ce^{4k}\] \[\frac{3}{2}=Ce^{2k}\] hmmm

OpenStudy (anonymous):

Correct.

OpenStudy (amistre64):

\[ln(5)-ln(C)=4k\] \[ln(3/2)-ln(C)=2k\] \[ln(5)-ln(C)=4k\]\[-2ln(3/2)+2ln(C)=-4k\] \[ln(5)-ln(C)-2ln(3/2)+2ln(C)=0\] \[ln(C)=ln(9/4)-ln(5)\] \[ln(C)=ln(9/20)\] \[C=9/20\] maybe?

OpenStudy (anonymous):

@amistre64, Can I try?

OpenStudy (amistre64):

\[ln(5)-ln(9/20)=4k\] \[ln(100/9)=4k\] \[\frac{1}{4}ln(100/9)=k\] its prolly wrong, but we can test it out at least :)

OpenStudy (amistre64):

yep, i was never good at the exponential finding stuff to well ;)

OpenStudy (anonymous):

What I got was: \[k = \ln (10/3) \times (1/2) \] and C = .45 but I am not sure those are correct.

OpenStudy (anonymous):

Divide to get K: e^2k = 10/3 2k = ln ( 10/3) ->k = (1/2) ln ( 10/3)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D%289%2F20%29+e%5E%28ln%28100%2F9%292%2F4%29 i think the wolf likes it

OpenStudy (anonymous):

Then plug it back: C = 5/ e^( 2ln10/3)

OpenStudy (amistre64):

\[\large y=\frac{9}{20}e^{\frac{1}{4}ln(100/9)t}\] works for me ... but like i said, i gots no idea if its what your looking for

OpenStudy (anonymous):

My PC and connection, including me are so slow! so pls do ask me if you need more explanation!

OpenStudy (amistre64):

9/20 = .45 so we agree on that http://www.wolframalpha.com/input/?i=ln%28100%2F9%29%2F4 says the 1/2 ln(10/3) is the same

OpenStudy (anonymous):

I just checked. \[y=.45e ^{\ln (10/3)\times(1/2)t} \] works with the points

OpenStudy (amistre64):

yep, it does :)

OpenStudy (anonymous):

Sweet! Thanks for helping!